OpenStudy (anonymous):

Solve for y in the equation 4y + 15 = –5y + 9. A. –6 B. 2003-03-04-00-00_files/i0280000.jpg C. 2003-03-04-00-00_files/i0280001.jpg D. 6

OpenStudy (anonymous):

4y+15=-5y+9 move -5y to the left side 4y+5y+15=9 subtract -15 from both sides 9y+15-15=9-15 9y=-6 y=-6/9 y=-2/3